You send out three messages for help:
a smoke signal that has prob .1 of being seen
a message in a bottle that has prob .2 of being found
a carrier pigeon that has prob .3 of arriving
Find the prob you are saved, assuming that smoke, bottles, and pigeons are independent.

Answer by ikleyn(53763)

.

It is a nice problem;  therefore,  I edited your post to make your formulation more clear.

    A shipwreck threw you up away to inhabitant island.
    You send out three messages for help:
        - a smoke signal that has the probability of 0.1 of being seen;
        - a message in a bottle that has the probability of 0.2 of being found, and
        - a carrier pigeon that has the probability of 0.3 of arriving.
    Find the probability your message will reach people and you will be saved, assuming 
    that these three signaling methods, smoke, bottles, and pigeons are independent.


Solution

Let's consider the complement event, that your message will not reach a resque crew.

The probability of this event is the product 


    P = (1-0.1)*(1-0.2)*(1-0.3) = 0.9*0.8*0.7 = 0.504.


Therefore, the probability that your message will reach people is the complement to it, i.e.


    P' = 1 - 0.504 = 0.496.      ANSWER

Solved.